OpenStudy (anonymous):

Give an example of an even function and explain algebraically why it is even.

Parth (parthkohli):

So, an even function \(f\) is a function where for all \(x\), \(f(x) = f(-x)\). The square function and the absolute-value functions are pretty cool examples.

OpenStudy (anonymous):

is that it?

Parth (parthkohli):

Maybe.

Parth (parthkohli):

Do you know what \(|x|\) is?

OpenStudy (anonymous):

x

Parth (parthkohli):

And \(|-x|\).

OpenStudy (anonymous):

x

Parth (parthkohli):

Bang on!

Parth (parthkohli):

And do you know what \((-x)^2\) is?

OpenStudy (anonymous):

x?

Parth (parthkohli):

Nope

OpenStudy (anonymous):

o

OpenStudy (anonymous):

x^2

Parth (parthkohli):

Yes.

Parth (parthkohli):

That's correct! =)

OpenStudy (anonymous):

so how does that answer this question, bro? :)

Parth (parthkohli):

You just answered the question, “explain algebraically why it is even”. Look back at the definition.

OpenStudy (hba):

If f(-a)=f(a) then it is even example f(x)=x^2 f(-x)=(-x)^2 f(-x)=x^2 therfore, f(x)=f(-x) @parth explained it well :)

OpenStudy (hba):

@ParthKohli *

Parth (parthkohli):

:)

Parth (parthkohli):

Let me do that for an absolute-value function too! So we define the absolute value function as follows\[f(x) = \cases{x \ \ \text{iff} \ \ x >0 \\ -x \ \ \text{iff} \ \ x < 0 }\]It is clear that \(f(-x)=f(x)=x\), hence an even function.